OpenStudy (anonymous):

Joey’s sock drawer is unorganized and contains 7 black dress socks, 7 black ankle socks, 6 brown dress socks, and 2 brown ankle socks. What is the probability that Joey will blindly reach into his sock drawer and pull out a sock that is brown or a dress sock? 4 d.

OpenStudy (anonymous):

number of brown or dress socks divided by total number of socks

OpenStudy (anonymous):

probability of brown sock = 8/22 probability of dress sock = 13/22 can u continue from here?

OpenStudy (anonymous):

So 21/22?